Interpretive Dimensions provides sign language interpreters for major business conferences. Interpreters are paid an hourly rate of $20. Fringe benefits are estimated to be 30 percent of wages, and employment taxes amount to 10 percent of wages.

Calculate the standard hourly rate for an interpreter. (Round answer to 2 decimal places, e.g. 52.75.)

Standard rate per hour $_____
Answer #1

Standard rate per hour

= 20 + 30% + 20%

= 20 + (20*30%) + (20*10%)

= 20 + 6 + 2

= 28 per hour